# Application of AI-based Models for Online Fraud Detection and Analysis

### v1

@@ 4.1 Data Sources
We find that many works collect user profile data from various online social networks for this analysis, like Twitter raj2020multi ; gangan2023detection ; yue2019madafe ; singh2023safe ; ali2019detect ; shukla2022tweezbot ; rovito2022evolutionary , Instagram das2022effecient ; fathima2023ann ; anklesaria2021survey , Facebook albayati2019identifying ; venkatesan2019graph ; shreya2022identification , YouTube na2023evolving , and Sina Weibo zhang2021social . A different approach was used by a study haq2023spammy that collected real names from various webpages, schools, and other sources to automatically detect fake names online.
@@
Other works used previously published and openly accessible datasets that included user data from various social networks nikhitha2023fake ; bebensee2021leveraging
@@
GenAI-facilitated Social Engineering Attacks. Under this category, we find many works that investigate how generative AI models can be misused to defraud people.
@@
Various studies janjeva2023rapid ; schmitt2023digital ; ferrara2024genai develop and discuss an initial taxonomy where they present how AI-generated content can be misused by scammers. At the same time, Carlini et al. carlini2021extracting test various membership inference attacks on OpenAI’s GPT2 model and confirm that the model is vulnerable to this kind of attack which poses risks of privacy. Similarly, Kumar et al. kumar2023augmenting discuss the significant implications for cybersecurity, privacy, and ethical considerations that should be considered when developing and using these models.
@@
Moving on to misuse cases of these models, ayoobi2023looming discuss how LLMs and GenAI can be used for fake professional profile bios to trick users that the account is legitimate. Similarly, diresta2024spammers show that scammers can use these models to create AI-generated images to be posted on various social networks. Their case study shows that these images tend to receive high volumes of engagement on Facebook as many users do not seem to recognize that the images are synthetic. Other works show how these models can be jailbroken to produce code to immitate legitimate webpages (phishing URLs) grbic2023social , malware code, phishing emails, phishing SMS, SQL injection attacks, and many more potentially dangerous material shibli2024abusegpt ; alotaibi2024cyberattacks ; alawida2024unveiling .
@@
At the same time other works show that humans may be able to accurately detect phishing AI-generated content sharma2023well , while Roy et al. roy2023chatbots discuss and experiment with countermeasure to prevent malicious prompts (jailbreaking) in GPT and they provide valuable insights into how the model can become more robust against this vulnerability.
@@
Social Media Scams. Six studies examine various scams and spammers facilitated through various Social Networks. xu2022efficiently used data from WeChat and Konect repository to detect users that use WeChat to defraud people. la2023sa and la2021uncovering used Telegram data to characterize and detect Fake Telegram channels, while  shah2020illicit collected data from Telegram and compared it to Twitter data to understand and detect fake users. Similarly, al2023dspamonto collect and analyzed Twitter and Institute of Informatics and Telematics data to detect scammers on Twitter. tripathi2022analyzing collect Youtube data to detect scammers that attempt to lure in victims via YouTube comments on videos.
GenAI Scam Baiting. We identified three works that used LLMs and Generative AI to automatically engage with scammers online towards wasting their resources and collect data on various fraud activities. For this section we do not identify data sources used as researchers use data received on their baiting accounts.
@@
Romance Fraud. Moving on to Romance Fraud, he2021datingsec attempt to automatically detect malicious accounts on Momo, a dating website. Similarly, suarez2019automatically collect data from DatingMore and scamdigger.com to develop automated methods to understand fraudulent profiles within dating social networks. Lastly, lokanan2023tinder analyzed the sentiment of tweets with the hashtag #tinderswindler to provide an understanding on users sharing their experiences regarding romance fraud.
Fraudulent Investment. Studies that attempted to understand fraudulent investment scams did so employing various datasets and methodologies.
@@
First, siu2022invest analyze investment scam advertisements found in Bitcointalk. [^26] https://bitcointalk.org/ Li et al. li2023towards collect YouTube comments to automatically detect bots that advertise fraudulent investment content. Lastly, kuo2024constructing develop a scam detection model based on emotional fluctuations of user discussions collected from one of the most popular instant messaging app in Taiwan.
Crypto Manipulation. Market, and more specifically, cryptocurrency coin manipulation is when users collectively attempt to alter investor interactions towards manipulating the prize of a coin.
nizzoli2020charting discuss this process in detail via data acquired from Twitter, Telegram, and Discord channels. Similarly, mirtaheri2021identifying identify and analyze cryptocurrency manipulations from user activity collected from Telegram and Twitter.
@@
Fraudulent Crowdfunding. Both works that analyse fraudulent crowdfunding lee2022backers ; shafqat2019topic collect the descriptions and metadata from hundreds of Kickstarter campaigns. [^27] https://www.kickstarter.com/
@@
Fraudulent E-commerce. salihovic2019role analyze the Terms and Conditions of various websites that sell various products towards understanding and detecting obscured financial obligations in online agreement.
@@ 4.2 Methodologies employed
We now discuss the most popular AI and NLP methodologies employed per online frayd type.
@@
Phishing URLs. We find that the records included in our SLR attempt to automatically detect phishing URLs using a variety of NLP and AI methodologies. These models included classic supervised machine learning algorithms such as Naive Bayes (NB), Random Forest (RF), Decision Tree (DT), Support Vector Machines (SVM), XGBoost (Extreme Gradient Boosting), KNN (k-Nearest Neighbors), as well as Artificial Neural Networks (ANN) and more advanced in deep learning such as Long Short-Term Memory (LSTM) and Convolutional Neural Network (CNN). LSTMs are a type of Recurrent Neural Network (RNN) designed to capture long-dependencies in sequential data, making them suitable for handling and predicting sequences of text. On the other hand, CNNs aim to identify key features in the text by capturing local patterns. These models were developed for binary classification tasks, whether a URL is fraudulent or not.
@@
NLP techniques related to text mining have been used to extract features from URLs, which are then used as features to train the AI-based classification models. More specifically, counts of characters, special characters, n-grams, and URL deconstruction: detection of secure scheme or not, domain and top-level domain, along with sub-directories. Various works used hybrid, or combination of methodologies for their detection including some more advanced techniques, like li2022mui using Bi-LSTM (Bidirectional Long Short-Term Memory that could process sequences of text in both forward and backward directions) with VGG (Visual Geometry Group - a type of CNN), vo2023shark using W2V (Word2Vec) for feature extraction along with CNN, GRU (Gated Recurrent Unit, - like LSTM but a more simplified RNN), and Bi-LSTM, nakano2023canary using BERT with RF, bitaab2023beyond developed a hybrid system that uses RF, SVM, FNN, and XGBoost, alswailem2019detecting used Linear Regression (LR) and DT, and villanueva2022application used LSTM and GRU.
@@
Other stand-alone AI methodologies, like LightGBM (Light Gradient Boosting Machine - an ensemble learning technique designed for handling large datasets with large features), RF, NB, and ANN also seem to work well on detecting phishing URLs, but by far the one with the best performance seems to be RF.
@@
Phishing Emails. The methodologies used for phishing email recognition weigh more on the NLP analysis. The majority of studies used various NLP methodologies for feature extraction, including, but not limited to topic analysis (LDA, BERT, BERTLARGE), n-gram extraction (TF-IDF, BOW, Clustering, W2V), and sentiment analysis (VADER, WordNet).
@@
Then, the studies that also opted for automated detection of phishing emails employed LLM analysis, RF, NB, SVM, CatBoost, LSTM, RNN, and many more.
@@
Phishing SMS. Similar to phishing email detection, phishing SMS detection also relies on state-of-the-art NLP methodologies, including LLMs, LDA, BERT and W2V. The existing literature also used AI methodologies like LR, SVM, CNN, GNN, LSTM, NB, and KNN for automated detection.
@@
Phishing Phone Calls. Studies on Vishing used various AI-methods for automated detection. The majority of studies used transcript text data for their analysis, with the exception of djire2023evaluating who analyzed deepfake voice analysis. For that study, the authors found that RNN had the best performance.
@@
On text data, various NLP and AI methodologies were used, including but not limited to SVM, NB, LSTM, CNN, RF, BERT, W2V, LR, and KNN.
Phishing User Reports. We find that for the analysis of user reports of various phishing activities, the use of BERT, SMO, J48, NB, RF, XGBoost, D2V, Jaccard, NER, and TF-IDF were applied.
@@
Fake Reviews. The most popular NLP methodologies that apply on Fake Review detection lean heavily on Sentiment Detection methodologies including VADER and WordNet. Similar to previous fraud analysis, the AI methods applied include various Neural Network Models like CNN.
@@
Fraudulent Job Postings. The methodologies employed for the automated detection of Fraudulent Job Postings include stand-alone machine learning algoritunsm used for classification tasks including LR, SVM, KNN, RF, XBoost, and ANN, and deep learning models including Bi-LSTM.
Scam Baiting. One of the Scam Baiting studies,  cambiaso2023scamming used OpenAI’s ChatGPT to reply to scammer emails. Similarly, bajaj2023automatic experimented with OpenAI’s ChatGPT and DistillBERT to categorize various scam emails they received and provided a qualitative analysis of how well the two models perform. The other study, chen2023active set up a mailserver and used data from Scambaiter mailserver, Enron Email Dataset, and ScamLetters.Info while employing their own DistillBERT model to engage with scammers towards categorizing and analyzing the various emails.
@@
Romance Fraud. Studies on Romance Fraud detection used various NLP methodologies, including sentiment detection using BOW, TF-IDF, and textBlob, finding that RF works best. For the detection of malicious accounts in dating applications, LSTM works best, while another study  suarez2019automatically found that EML also works well.
@@
Fraudulent Investment. Three of the studies related to Fraudulent Investment used data from different sources (forum, messaging app, and YouTube). The reported studies indicate that DT was the best performed machine learning model for detecting emotional flactuations on victims, while XGBoost worked well on detecting malicious advertisement of fraudulent investment websites.
Crypto Manipulation. The two identified studies on Cryptocurrency Market Manipulation used pre-existing methods for detecting fake users, along with CorEx Topic Analysis nizzoli2020charting . The other study found that SVM with SGD and TF-IDF works best for detecting discussions that aim to manipulate the market mirtaheri2021identifying .
@@
Fraudulent Crowdfunding. One of the papers on fraudulent crowdfunding detection used NLP methodologies including Named Entity Recognition and other NLP featurues detected in the descriptions of Kickstarter campaigns and build a LR model that performs well lee2022backers . The other study, shafqat2019topic developed an LSTM-LDA topic detection model that analyses the crowdfunding campaign, along with the comments of people towards estimating whether a campaign is a scam or not.
@@
Fraudulent E-Commerce. Finally, the only study that we identified that used text data towards understanding Fraudulent E-Commerce activities, used OpenAI’s GPT-4 model to automatically detect obscure financial obligations in the Terms and Conditions of the focused websites salihovic2019role .

### v2

@@ 5.1 Data Sources
We found that many authors collected user profile data from online social networks including Twitter raj2020multi ; gangan2023detection ; yue2019madafe ; singh2023safe ; ali2019detect ; shukla2022tweezbot ; rovito2022evolutionary , Instagram das2022effecient ; fathima2023ann ; anklesaria2021survey , Facebook albayati2019identifying ; venkatesan2019graph ; shreya2022identification , YouTube na2023evolving , and Sina Weibo zhang2021social . A different approach used in one study haq2023spammy was to collect real names from various web pages, schools, and other sources to detect fake names online automatically.
@@
Other authors have used previously published and openly accessible datasets that included user data from various social networks nikhitha2023fake ; bebensee2021leveraging .
@@
Data Used for GenAI Social Engineering Attack Detection. Under this category, we found many works that investigated how generative AI models can be misused to defraud people.
@@
Some studies janjeva2023rapid ; schmitt2023digital ; ferrara2024genai develop and discuss an initial taxonomy for which they discuss how scammers can misuse AI-generated content. At the same time, Carlini et al. carlini2021extracting test various membership inference attacks – which is when someone attempts to figure out whether a specific piece of data was used to train a machine learning model – on OpenAI’s GPT2 model and confirm that the model is vulnerable to this kind of attack which poses risks to privacy. Similarly, Kumar et al. kumar2023augmenting discuss the significant implications for cybersecurity, privacy, and ethical considerations that should be considered when developing and using these models.
@@
Apropos misuse cases of these models, Ayoobi et al. ayoobi2023looming discuss how LLMs and GenAI can be used to create fake professional profile bios to trick users into believing that the account is legitimate. Similarly, DiResta and Goldstein diresta2024spammers show that scammers can use these models to create AI-generated images to be posted on social networks. Their case study shows that these images tend to receive high volumes of engagement on Facebook as many users do not seem to recognize that the images are synthetic. Other research shows how these models can be jailbroken [^26] Jailbreaking a generative AI model means bypassing its safety rules or restrictions to make it produce responses it’s not supposed to. to produce code to imitate legitimate webpages (phishing URLs) grbic2023social , malware code, phishing emails, phishing SMSs, SQL injection attacks, and other potentially malicious material shibli2024abusegpt ; alotaibi2024cyberattacks ; alawida2024unveiling .
@@
Other research suggests that humans may be able to accurately detect phishing AI-generated content sharma2023well , while Roy et al. roy2023chatbots discuss and experiment with countermeasures to prevent malicious prompts (jailbreaking) for GPT and provide insights into how the model can become more robust against this vulnerability.
@@
Data Used for Social Media Scam Detection. Six studies examined various scams and spammers facilitated by Social Networks. Xu et al. xu2022efficiently used data from WeChat (a Chinese messaging, social media, and mobile payment app) and Konect repository to detect users that use WeChat to defraud people. La Morgia et al. (2023) la2023sa and La Morgia et al. (2021)  la2021uncovering used Telegram data to characterize and detect Fake Telegram channels, while Shah et al. shah2020illicit collected data from Telegram and compared it to Twitter data to understand and detect fake users. Similarly, Al-Hassan et al. al2023dspamonto collected and analysed Twitter and Institute of Informatics and Telematics data to detect scammers on Twitter. Finally, Tripathi et al. tripathi2022analyzing collected YouTube data to detect scammers attempting to lure victims using comments posted alongside YouTube videos.
@@
Data Used for Romance Fraud Detection. In their study, He et al. he2021datingsec attempted to automatically detect malicious accounts on Momo, [^27] https://www.immomo.com/aboutus.html a dating website. Similarly, Suarez-Tangil et al. suarez2019automatically collected data from datingnmore.com and scamdigger.com to develop automated methods to understand fraudulent profiles within dating social networks. Lastly, Lokanan lokanan2023tinder analysed the sentiment of tweets with the hashtag #tinderswindler to provide an understanding of users sharing their experiences regarding romance fraud.
Data Used for GenAI Scam Baiting. We identified three studies. cambiaso2023scamming ; bajaj2023automatic ; chen2023active that used LLMs and Generative AI to automatically engage with scammers online to waste their resources and collect data on various fraud activities. For these studies, the researchers collected data from their own baiting accounts and emails and said data was not made publicly available.
@@
Data Used for Fraudulent Investment Detection. Studies that attempted to understand fraudulent investment scams employed various datasets and methodologies. First, Siu et al. siu2022invest analyse investment scam advertisements found in Bitcointalk. [^28] https://bitcointalk.org/ Li et al. collected YouTube comments to detect bots that advertise fraudulent investment content automatically li2023towards . Lastly, Kuo and Tsang  kuo2024constructing develop a scam detection model based on emotional fluctuations of user discussions collected from one of Taiwan’s most popular instant messaging applications.
@@
Data Used for Fraudulent Crowdfunding Detection. Two studies were identified that analyse fraudulent crowdfunding lee2022backers ; shafqat2019topic . These collected the descriptions and metadata from hundreds of Kickstarter campaigns. [^29] https://www.kickstarter.com/
Data Used for Crypto Manipulation Detection. Market, and more specifically, cryptocurrency coin manipulation, is when users collectively attempt to alter investor interactions towards manipulating the price of a coin.
Nizzoli et al. nizzoli2020charting discuss this process via data acquired from Twitter, Telegram, and Discord channels. Similarly, Mirtaheri et al. mirtaheri2021identifying identify and analyse cryptocurrency manipulations from user activity collected from Telegram and Twitter.
@@
Data Used for Fraudulent E-commerce Detection. The only study for this category salihovic2019role analysed the terms and conditions of websites that sell (a variety of) products to inform understanding and the detection of obscured financial obligations in online agreements.
@@ 5.2 Methodologies employed
We now discuss the most popular AI and NLP methodologies employed to study each type of online fraud.
@@
Methods Applied for Phishing URL Detection. The studies included in our SLR attempted to automatically detect phishing URLs using a variety of NLP and AI methodologies. These included classic supervised machine learning algorithms such as Naive Bayes (NB), Random Forest (RF), Decision Tree (DT), Support Vector Machines (SVM), XGBoost (Extreme Gradient Boosting), KNN (k-Nearest Neighbors), as well as Artificial Neural Networks (ANN) and more advanced deep learning approaches such as Long Short-Term Memory (LSTM) and Convolutional Neural Network (CNN). LSTMs are Recurrent Neural Networks (RNN) designed to capture long-dependencies in sequential data, making them suitable for handling and predicting text sequences. On the other hand, CNNs aim to identify key features in the text by capturing local patterns. In the research reviewed, these models were developed to complete the binary classification task of determining whether a URL was fraudulent or not.
@@
NLP techniques related to text mining have been used to extract features from URLs, which are then used as features to train an AI-based classification model. For example, such features include the counts of the characters, special characters, and n-grams of the URLs. Also, the authors collected other URL features, such as whether the URL had a secure scheme or not (e.g. https), domain (e.g. amazon.co.uk) and top-level domain (e.g. /kitchen), and sub-directories (e.g. /appliances). All the above and more features were used to fit and train a malicious URL detection model. Some studies used hybrid or a combination of methodologies, including more advanced techniques. For example, Li et al. li2022mui used a Bidirectional Long Short-Term Memory (Bi-LSTM) recurrent neural network, that could process sequences of text in both forward and backward directions, along with a Visual Geometry Group (VGG) which is a type of CNN. Vo Quang et al. vo2023shark used a Convolutional Neural Network (CNN), along with features extracted using Word2Vec (W2V), a Gated Recurrent Unit (GRU), which is a more simplified type of RNN than LSTM, and a Bi-LSTM. Nakano et al. nakano2023canary used BERT with RF; Bitaab et al. bitaab2023beyond developed a hybrid system that uses RF, SVM, FNN, and XGBoost; Alswailem et al. alswailem2019detecting used Linear Regression (LR) and DT; and, Villanueva et al. villanueva2022application used LSTM and GRU.
@@
Other stand-alone AI methodologies, like LightGBM, [^30] Light Gradient Boosting Machine - an ensemble learning technique designed for handling large datasets with large features RF, NB, and ANN, also seem to work well on detecting phishing URLs, but the approach with the best performance seems to be RF.
@@
Methods Applied for Phishing Email Detection. The methodologies used for phishing email recognition focus more on NLP analysis. The majority of studies used various NLP methodologies for feature extraction, including, but not limited to topic modeling (LDA, BERT, BERTLARGE), text representation (TF-IDF, BOW, Clustering, W2V), and sentiment analysis (VADER, WordNet).
@@
Studies that also aimed to automate the detection of phishing emails employed LLM analysis, RF, NB, SVM, CatBoost, LSTM, RNN, and many more.
@@
Methods Applied for Phishing SMS Detection. Like phishing email detection, phishing SMS detection relies on state-of-the-art NLP methodologies, including LLMs, LDA, BERT and W2V. The existing literature also used AI methodologies like LR, SVM, CNN, GNN, LSTM, NB, and KNN for automated detection.
@@
Methods Applied for Phishing Phone Call Detection. Studies on vishing detection used various AI methods for automated detection. Most used transcript text data for their analysis, except for Djiré et al. djire2023evaluating who analysed deepfake voice analysis. In that study, the authors found that RNNs performed best.
@@
Overall, various NLP and AI techniques were used on text data, including but not limited to SVM, NB, LSTM, CNN, RF, BERT, W2V, LR, and KNN.
Methods Applied for Phishing (User Reports) Detection. We found that the use of BERT, Sequential Minimal Optimization (SMO), J48 (an implementation of decision tree), NB, RF, XGBoost, Doc2Vec (D2v - an extension of Word2Vec), Jaccard, NER (Named Entity Recognition), and TF-IDF was applied to analyse user reports of various phishing activities.
@@
Methods Applied for Fake Review Detection. The most popular NLP technique for fake review detection rely heavily on sentiment detection techniques, including VADER and WordNet. Similar to previous fraud analyses, the AI methods applied included various neural network models, such as CNN.
@@
Methods Applied for Recruitment Fraud Detection. The techniques employed for the automated detection of Fraudulent Job Postings included stand-alone machine learning algorithms used for classification tasks, including LR, SVM, KNN, RF, XBoost, and ANN, and deep learning models, including Bi-LSTM.
Methods Applied for Fake Account Detection. The studies included in our SLR leveraged various NLP and AI-based techniques to detect fake accounts on social media platforms such as Twitter, Facebook, Instagram, and YouTube. Classic supervised learning approaches, including NB, RF, DT, SVM, LR, KNN, and ANN, were widely adopted. Ensemble learning methods such as AdaBoost and stacking models were also explored, along with advanced methods like Gradient Boosting techniques (e.g., CatBoost). Notably, RF was often found to deliver the best performance in several studies, e.g., anklesaria2021survey ; nikhitha2023fake ; das2022effecient .
Deep learning approaches were also employed, particularly when tackling larger datasets. For example, Na et al. na2023evolving used RoBERTa to detect fake accounts involved in scam campaigns on YouTube, while Alhosseini et al. ali2019detect leveraged Graph Convolutional Neural Networks (GCNN) for spam bot detection on Twitter. Studies such as venkatesan2019graph adopted unsupervised learning techniques like HDBSCAN for anomaly detection in social networks.
Across the studies reviewed, researchers have extracted diverse features, including user profile characteristics (e.g., number of followers, account age), content-based features (e.g., hashtags, posts), domain-based features, and behavioural patterns.
Deep learning models such as Bi-LSTM, GRU, and CNN were less frequently applied but showed promise. Fathima et al. fathima2023ann developed an ANN-based system to categorise fake profiles on Instagram. Ensemble learning approaches, such as combining ANN, SVM, and RF shreya2022identification , were also utilised to enhance detection performance.
The performance of these methods varied depending on the dataset and feature selection, but overall, RF emerged as the most consistent and accurate classifier for fake account detection across multiple platforms and studies.
Methods Applied for GenAI Social Engineering Attack Detection. GPT-3.5 and GPT-4 were the most commonly used models, particularly in generating phishing emails, malicious websites, and smishing campaigns. Studies by Roy et al. roy2023chatbots and Shibli et al. shibli2024abusegpt demonstrated how these models could be exploited to craft highly convincing phishing content, leveraging sophisticated language capabilities. Ayoobi et al. ayoobi2023looming , utilised models like BERT, RoBERTa, and Flair to detect fake LinkedIn profiles generated by ChatGPT. Defensive mechanisms were also explored; for instance, Roy et al. roy2023chatbots proposed BERT-based countermeasures to mitigate malicious prompt exploitation. Despite the promising results in detecting and preventing misuse, other research  alotaibi2024cyberattacks  alawida2024unveiling has highlighted vulnerabilities in GPT-3.5, particularly its susceptibility to jailbreak attempts, which enable the generation of harmful content such as SQL injections, malware, and phishing scams. Overall, LLMs demonstrated advanced capabilities for deception, and their susceptibility to misuse necessitates robust detection and prevention strategies. Studies under this category did not report performance metrics as they tested the limits of LLMs and GenAI models using qualitative approaches.
Methods Applied for Social Media Scam Detection. Xu et al. xu2022efficiently proposed the BREAD framework, which uses bidirectional k-hop reachability query processing over dynamic graphs to extract fraud-related features. La Morgia et al. (2023) la2023sa studied fake Telegram channels employing a Multilayer Perceptron (MLP) model. Shah et al. shah2020illicit applied techniques like W2V, D2V, P2V, and TF-IDF to detect illicit activity. Tripathi et al. tripathi2022analyzing examined monetised scam videos on YouTube using RF and W2V. Al-Hassan et al. al2023dspamonto developed DSpamOnto, an ontology-based model for social spammers on Twitter, and benchmarked it against classifiers such as NB, SVM, and RF. Finally, La Morgia et al. (2021) la2021uncovering used LDA for topic modelling.
Due to the diversity in the types of fraud analysed across social media platforms and the distinct datasets and methodologies employed, identifying a single best-performing model for this category is not applicable.
@@
Methods Applied for Romance Fraud Detection. Studies of Romance Fraud detection used various NLP methodologies, including sentiment detection, which uses BOW and textBlob, and statistical methods like TF-IDF, for feature extraction. When testing different models, researchers have found that Random Forest (RF) performed best for the detection of this offence lokanan2023tinder . He et al. he2021datingsec found that LSTM is most effective at detecting malicious accounts in dating applications, while another study suarez2019automatically showed that Ensemble Machine Learning (EML) also works well.
Methods Applied for GenAI Scam Baiting. One of the Scam Baiting studies cambiaso2023scamming used OpenAI’s ChatGPT to reply to scammer emails. Similarly, Bajaj and Edwards bajaj2023automatic experimented with OpenAI’s ChatGPT and DistillBERT to categorize scam emails they received and provided a qualitative analysis of how well the two models performed. Chen et al. chen2023active set up an email server as a “honeypot” from which they sent emails to scammers (to encourage those scammers to send phishing emails to them) identified in data from the Scambaiter mail server, Enron Email Dataset, and ScamLetters.Info. They then employed their own semi-unsupervised DistillBERT model to engage with scammers automatically and used their model filtering to categorise and analyse the emails received.
@@
Methods Applied for Fraudulent Investment Detection. Three of the studies related to Fraudulent Investment used data from different sources (forums, messaging apps, and YouTube). Two of the studies kuo2024constructing ; siu2022invest reviewed under this SLR used models to detect emotional fluctuations in discussions between victims and found that DT was the best-performing machine learning model for this task. Siu et al. siu2022invest also concluded that XGBoost performed well in terms of detecting malicious advertisements for fraudulent investment websites.
@@
Methods Applied for Fraudulent Crowdfunding Detection. One of the papers on fraudulent crowdfunding detection applied NLP methodologies, including Named Entity Recognition and other NLP features detected in the descriptions of Kickstarter campaigns, and built an LR model that performed well lee2022backers . The other study shafqat2019topic developed an LSTM-LDA topic detection model that analyses the crowdfunding campaign and people’s comments with the aim of estimating whether a campaign was a scam.
Methods Applied for Crypto Manipulation Detection. One study on cryptocurrency market manipulation used pre-existing methods for detecting fake users, along with CorEx Topic Analysis nizzoli2020charting . The other study found that SVM with SGD and TF-IDF worked best for detecting discussions that aimed to manipulate the market mirtaheri2021identifying .
@@
Methods Applied for Fraudulent E-Commerce Detection. Finally, the only study that we identified that used text data to inform understanding of Fraudulent E-commerce activities, used OpenAI’s GPT-4 model to automatically detect obscure financial obligations in the terms and conditions of the websites sampled salihovic2019role .